OpenStudy (anonymous):

mathematical induction to prove that n 3 − n is divisible by 3, for every positive integer n

OpenStudy (solomonzelman):

\(n^3-n=n(n^2-1)=n(n-1)(n+1)\) that should help.

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

No further questions?

OpenStudy (anonymous):

not yet... im working it thru

OpenStudy (anonymous):

ok this is where im at P(x+1)=(x+1)((x+1)−1)((x+1)+1) = 3m do i divide

OpenStudy (solomonzelman):

you don't need anything, except for a little logic.

OpenStudy (solomonzelman):

here, tell me what happens if you have a product that consists of integers and one of these integers are divisible by 3? Do you agree that the result is divisible by 3?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, good.

OpenStudy (solomonzelman):

Now, lets come back to the fact that you want to prove that: *n(n-1)(n+1)* is divislbe by 3, \(\forall {\bf n \in \mathbb Z}\)

OpenStudy (solomonzelman):

Ok, lets consider 3 possible cases (for possible integer k) 3k, 3k+1, and 3k+2

OpenStudy (solomonzelman):

If your number n falls under the category 3k (Such that n --> 3k) then the *n* component of *n(n-1)(n+1)*, is divisible by 3, and thus the entire product *n(n-1)(n+1)* is divislbe by 3.

OpenStudy (solomonzelman):

If your number n falls under the category 3k+1 (Such that n --> 3k+1) then the *n-1* component of *n(n-1)(n+1)*, is divisible by 3, and thus the entire product *n(n-1)(n+1)* is divislbe by 3.

OpenStudy (solomonzelman):

And then if: n --> 3k+2 then the *n+1* makes it divisible by 3.

OpenStudy (solomonzelman):

Do I sound rediculous, or is it understandable.

OpenStudy (anonymous):

no u dont.. im just trying to absorb the concept

OpenStudy (anonymous):

my brain is a bit mathed out

OpenStudy (solomonzelman):

\(\displaystyle\int\)\(\theta\) \(f(u)+nn=y\)

OpenStudy (solomonzelman):

If you have more questions to ask, I wil be back if I am online.

OpenStudy (anonymous):

thank you